The magnitude of the electrostatic force between two identical ions that are separated by a distance of 6.6 x 1010 m is 331.1 x 10-9 N. (a) What is the charge of each ion? (b) How many electrons are missing from each ion (thus giving the ion its charge imbalance)? (a) NumberT 4.005E-19 (b) NumberT 2.5 Units C - u Units No units
Answer #1

a)

F = K*q*q/d^2

F = K*q^2/d^2

331.1*10^-9 = (9*10^9)*(q)^2 / (6.6*10^-10)^2

q^2 = 1.602*10^-35

q = 4.00*10^-18 C

Answer: 4.00*10^-18 C

b)

charge of 1 electron = 1.602*10^-19 C

So,

number of electrons = q / charge of 1 electron

= (4.00*10^-18 C) / (1.602*10^-19 C)

= 25.0

Answer: 25